OpenStudy (anonymous):

six divided by a number is 1/2

OpenStudy (anonymous):

6/x = 1/2 6 = x/2 x = 12

OpenStudy (anonymous):

i put answer of 12 and it was counted wrong

OpenStudy (anonymous):

it is correnct.

OpenStudy (anonymous):

ok thank you

myininaya (myininaya):

it says to solve for that number right?

OpenStudy (anonymous):

6/x = 12.. dont solve for x

OpenStudy (anonymous):

the exact question is 6 divided by a number is 1/2

myininaya (myininaya):

that doesn't say to do anything lol

myininaya (myininaya):

that is just a statement

OpenStudy (anonymous):

i'm typing it right out of the book they are writing expressions and equations

OpenStudy (anonymous):

he is converting word problems into actual equations is how im interpreting it

myininaya (myininaya):

so no where in that section of answering questions does it to solve?

OpenStudy (anonymous):

ok, so it will be, 6/x = 1/2

OpenStudy (anonymous):

writing verbal phrase or sentence as an algebraic expression or equation

OpenStudy (anonymous):

the answer is \[\frac{6}{x} = 12\]

myininaya (myininaya):

ok so yes 6/x=1/2 would be write

myininaya (myininaya):

right*

OpenStudy (anonymous):

the other kids are correction papers in the class not the teacher

myininaya (myininaya):

what?

OpenStudy (anonymous):

she had the right answer and they marked her wrong

OpenStudy (anonymous):

i just noticed i wrote 12 instead of 1/2.. its suppose to be 6/x = 1/2